OpenStudy (savvygrl108):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (supersmart1001):

1. sin2x = -cos(-x + 9) = cos(-x +9 +180) 2. but 2x - x + 9 + 180 = 90 3. x = -99

OpenStudy (supersmart1001):

thats all i got

OpenStudy (zpupster):

add 5 to both sides 7 sin2x - 14 sin x + 7 = 0 factor 7(sin^2x -2sinx + 1) = 0 divide both by 7 and factor do you see it now??

OpenStudy (karim728):